### v1

# Matchings and loose cycles in the semirandom hypergraph model

Abstract.
We study the 2-offer semirandom 3-uniform hypergraph model on $n$ vertices. At each step, we are presented with 2 uniformly random vertices. We choose any other vertex, thus creating a hyperedge of size 3. We show a strategy that constructs a perfect matching, and another that constructs a loose Hamilton cycle, both succeeding asymptotically almost surely within $\Theta(n)$ steps. Both results extend to $s$-uniform hypergraphs. Much of the analysis is done on an auxiliary graph that is a uniform $k$-out subgraph of a random bipartite graph, and this tool may be useful in other contexts.
1. Introduction and Main Results
1.1. The model
We consider the semirandom graph process suggested by Peleg Michaeli and introduced formally in [3]. This model, that can be viewed as a “one player game”, generalizes many well-known and important processes such as the Erdős-Rényi random graph process and the $k$-out process, and has received a lot of attention recently. We concentrate on its natural generalization to hypergraphs, as proposed in [1].
In the $r$-offer $s$-uniform semirandom model, the player constructs a sequence of $s$-uniform hypergraphs $H_{t}$ on vertex set $[n]:=\{1,\ldots,n\}$, with the goal that $H_{t}$ satisfies a given property $\mathcal{P}$ as quickly as possible. In each round $t\in\mathbb{N}$, a uniformly random set of $r$ vertices is offered, and the player chooses $s-r$ additional vertices according to some strategy, to make a hyperedge that is added to $H_{t-1}$ to form $H_{t}$. When choosing the $s-r$ vertices, the player has full knowledge of the $r$ vertices offered and of the past, including the hypergraph $H_{t-1}$ at the end of the previous round.
Results presented in this paper are asymptotic by nature. We say that a property $\mathcal{P}$ holds asymptotically almost surely (a.a.s.) if the probability that $H$ has property $\mathcal{P}$ tends to $1$ as $n$ goes to infinity.
In our first main result, we concentrate on the 2-offer semirandom construction of a 3-uniform hypergraph $H$ on $n$ vertices, $n$ divisible by 3, so that $H$ will contain a perfect matching (a partition of the set of vertices into $n/3$ hyperedges).
Theorem 1.
For some constant $C$, the 2-offer 3-uniform semirandom model has a strategy that, for $n$ divisible by $3$, in time $t=Cn$, gives a hypergraph $H_{t}$ a.a.s. containing a perfect matching.
We present the strategy explicitly, and show how to construct a matching in $H$ (not merely proving its existence). This result is clearly best possible up to the constant $C$, which we have made no effort to optimise.
Our linear-time construction extends to perfect matchings in an $s$-uniform hypergraph when the player is offered $r=2$ vertices and is able to select $s-2$ vertices. Moreover, it immediately implies the same result for a “1-offer” semirandom model, as the player can simply simulate a 2nd random offer vertex with the first of the $s-1$ chosen vertices.
Theorem 2.
For any $s\geq 3$ and $r\in\left\{1,2\right\}$, for some constant $C$, the $r$-offer $s$-uniform semirandom model has a strategy that, for $n$ divisible by $s$, in time $t=Cn$, gives a hypergraph $H_{t}$ a.a.s. containing a perfect matching.
Our second main result is a somewhat similar strategy to construct a 3-uniform hypergraph $H$ on $n$ vertices, $n$ divisible by 2, so that $H$ will contain contain a loose Hamilton cycle.
Theorem 3.
For some constant $C$, the 2-offer 3-uniform semirandom model has a strategy that, for $n$ divisible by $2$, in time $t=Cn$, gives a hypergraph $H_{t}$ a.a.s. containing a loose Hamilton cycle.
This result also extends.
Theorem 4.
For any $s\geq 3$ and $r\in\left\{1,2\right\}$, for some constant $C$, the $r$-offer $s$-uniform semirandom model has a strategy that, for $n$ divisible by $s-1$, in time $t=Cn$, gives a hypergraph $H_{t}$ a.a.s. containing a loose Hamilton cycle.
1.2. Outline
The next section gives some background on the semirandom process. Section 4 proves Theorem 1 by presenting a 2-offer strategy for constructing a 3-uniform hypergraph matching, and analysing it. Subsection 4.4 proves the extension to $s$-uniform hypergraphs. This is a simple modification of the argument for the 3-uniform case but requires reference to the strategy: Theorem 2 is not a black-box corollary of Theorem 1. Section 5 proves Theorems 3 and 4. The strategy for constructing a loose Hamilton cycle is different in its details from that for matching, but has the same structure and shares some key elements.
2. Background
We begin with a clarification. Recall that our $r$-offer, $s$-uniform semirandom hypergraph is formed as follows: At each round $t$, the player is offered $r$ uniform vertices and then chooses $s-r$ additional vertices to form a hyperedge. In our proofs and previous works, there are situations where, upon seeing the offered vertices, the player decides not to build any hyperedge on them and so does not bother choosing $s-r$ vertices. We say that the player ignores the offer. Formally, we can think of the player adding an arbitrary set of $s-r$ vertices and then never using the resulting hyperedge. This counts as a round and so, e.g., is one of the $Cn$ steps in Theorem 1.
Since the semirandom process is still relatively new, let us highlight a few results on this model, focusing first on perfect matchings.
In the very first paper [3], it was shown that the semirandom process on graphs is general enough to simulate several well-studied random graph models (using suitable strategies), including the extensively studied $k$-out process. In the $k$-out process on $n$ vertices, each vertex independently connects to $k$ randomly selected vertices, resulting in a random graph with $kn$ edges (see, for example, Chapter 18 in [16]). This is easily simulated with the semirandom process by “choosing” vertex 1 to accompany each of the first $k$ offered vertices, choosing vertex 2 for the next $k$, etc.
Since the $2$-out process has a perfect matching a.a.s. [9], we immediately get that a.a.s. one can create a perfect matching in $(2+o(1))n$ rounds. Using the semirandom process to construct a different sort of random graph known to have a perfect matching a.a.s. [17], the upper bound can be improved to $(1+2/e+o(1))n<1.73576n$ [3]. This bound was subsequently improved by investigating a fully adaptive algorithm. The current best upper bound is $1.20524n$ [13]. On the other hand, [3] observed that $(\ln(2)+o(1))n>0.69314n$ is a lower bound on the number of rounds needed to create a perfect matching, and this has since been improved to $0.93261n$ [13].
One approach to constructing perfect matchings in the 2-offer 3-uniform model would be to use a random $k$-out hypergraph, i.e., a hypergraph with $k$ random edges on (and associated with) each vertex. This is easily simulated in the semirandom process just as for graphs, choosing vertex 1 with each of the first $k$ offers, etc. It is natural to expect that such a hypergraph has a perfect matching, but it is only known to have a perfect fractional matching [5]. As said in [5], proving that it has a perfect matching is “likely to be difficult, as [it] would imply the … resolution of Shamir’s Problem, as well as [a] natural guess regarding a beautiful problem of Frieze and Sorkin [7].” Our approach is quite different.
Let us also briefly comment on other results on the semirandom graph model. Unsurprisingly, Hamilton cycles have received a lot of attention [3, 11, 12, 8, 6], and currently the best upper and lower bounds are $1.81701n$ and $1.26575n$ [6]. More general spanning subgraphs were considered in [2]. It showed that any graph with maximum degree $\Delta$ can be constructed a.a.s. in $(3\Delta/2+o(\Delta))n$ rounds; it also showed that if $\Delta=\omega(\log(n))$, this upper bound improves to $(\Delta/2+o(\Delta))n$ rounds, close to the trivial lower bound of $\Delta n/2$. Note, however, that these results are meaningful only for $\Delta\rightarrow\infty$; for any fixed $\Delta$ they say only that $O(n)$ rounds suffice. Cliques, chromatic number, and independent sets were considered in [10].
Finally, while our interest is in hypergraphs, let us mention some other variants of the semirandom process. In [18], sharp thresholds were studied for a more general class of processes that includes the semirandom process. In [4], a random spanning tree of $K_{n}$ is presented, and the player keeps one of the edges. In [14], vertices are presented by the process in a random permutation. In [20], the process presents $k$ random vertices, and to create an edge the player selects one of them, and freely chooses a second vertex.
3. Outline, and lemmas on an auxiliary bipartite graph
A uniformly random $k$-out bipartite multigraph (or just “uniform $k$-out” graph) on disjoint node sets $\mathcal{A},\mathcal{B}$ is a directed graph $\mathcal{D}$ in which every node in $\mathcal{A}$ has outedges to $k$ nodes in $\mathcal{B}$ chosen uniformly at random with replacement, and symmetrically for $\mathcal{B}$.
A uniform $k$-out auxiliary graph plays a central role in our semirandom construction. To motivate the lemmas in this section, we briefly explain the role each will play. Proper explanations will come in later sections.
Throughout, we will refer to vertices of the auxiliary graph as “nodes”, calling vertices of the semirandom process itself “points”, “elementary vertices”, or occasionally “vertices”. We will use calligraphic letters in this section, for general node sets $\mathcal{A}$ and $\mathcal{B}$ and a corresponding $k$-out graph $\mathcal{D}$. In context, later, we will define particular node sets $A$ and $B$ and a corresponding $k$-out graph $D$.
The number of nodes in the auxiliary graph is different from the number of points in the hypergraph. We will reserve $n$ for the latter, writing $N$ in the lemmas below to make it easy later to define $N$ in terms of $n$. The $k$ in the lemmas is the same as that in the main argument; indeed, we will take $k=10$ throughout.
Our semirandom constructions of both hypergraph matchings and loose Hamilton cycles will begin by defining node sets $\mathcal{A}$ and $\mathcal{B}$ (differently for the two cases).
Phase 1 will form a uniform $k$-out graph from a subset of a set of uniformly random edges on $\mathcal{A}\times\mathcal{B}$. Because the edges supplied are uniformly random, a few nodes are bound to have too few edges, but Lemma 6 shows that we can get a uniform $k$-out graph $\mathcal{D}$ on most of the nodes in $\mathcal{A}$ and $\mathcal{B}$. The few nodes missing from $\mathcal{D}$ will correspond to points of the hypergraph we will have to deal with specially.
Phase 2 will deal with these points. The “actions” addressing them will have the side effect of deleting more nodes from $\mathcal{D}$, correspondingly decreasing the degrees of the remaining nodes. We will arrange that the outdegree of a node in $\mathcal{D}$ is always either $k$ or $k-1$, by ensuring that each outdegree is reduced at most once (i.e., by one action), and by at most 1. The “by at most 1” comes from ensuring that we never delete a node with parallel inedges (7 shows that such nodes are rare), and that we never simultaneously delete a set of nodes sharing an inneighbour (8 shows that such sets are rare). (Alternatively, “by at most constant” is immediate from the nature of the actions used, and we could have used a larger $k$ to end with the same outdegree bound.) The “at most once” is achieved by “blocking” a deleted node’s inneighbour’s outneighbours from future deletion, adding these nodes to a set $Q$ of nodes not to be deleted; Lemma 9 is the basis for showing that the set $Q$ remains small. After all the actions, $\mathcal{D}$ still contains most of $\mathcal{A}$ and $\mathcal{B}$.
Phase 3 will complete the hypergraph construction using $\mathcal{D}$ as an absorber. Walkup [21] showed that a random 2-out bipartite graph has a perfect matching a.a.s., and Frieze [9] showed the same without the bipartiteness condition, for 2-out graphs and multigraphs. Here, despite the fact that $\mathcal{D}$ is neither exactly $k$-out nor uniformly random, Lemma 10 shows that almost certainly Hall’s condition is satisfied, so it has a perfect matching. The hypergraph construction will be completed by hyperedges corresponding to the matching edges.
Throughout, we rely on McDiarmid’s form of the Azuma-Hoeffding inequality [19].
Theorem 5 (Azuma).
Let $X_{1},\ldots X_{n}$ be independent random variables with $X_{k}$ taking values in a set $A_{k}$ for each $k$. Suppose that the (measurable) function $f\colon\prod A_{k}\rightarrow\mathbb{R}$ satisfies $\left|{f(x)-f^{\prime}(x)}\right|\leq c_{k}$ whenever the vectors $x$ and $x^{\prime}$ differ only in the $k$th coordinate. Let $Y$ be the random variable $f(X_{1},\ldots,X_{n})$. Then for any $t>0$, $\Pr(\left|{Y-\mathbb{E}Y}\right|\geq t)\leq 2\exp(-2t^{2}/\sum c_{k}^{2})$.
Lemma 6 ($k$-out).
For all integers $k>0$ and reals $\epsilon>0$, there exists a constant $C>0$ such that the following holds. Let $\mathcal{A}$ and $\mathcal{B}$ be disjoint nonempty sets with $\mathcal{A}\cup\mathcal{B}=[N]$.
Let $\Psi$ be a multiset of $CN$ (or more) uniformly random undirected edges on $\mathcal{A}\times\mathcal{B}$. Then, a.a.s., we can delete some edges from $\Psi$, and orient the remaining ones, to give a directed bipartite multigraph $\mathcal{D}$ on parts $\mathcal{A}^{\prime}\subseteq\mathcal{A}$ and $\mathcal{B}^{\prime}\subseteq\mathcal{B}$, where $\left|{\mathcal{A}^{\prime}}\right|\geq(1-\epsilon)\left|{\mathcal{A}}\right|$, $\left|{\mathcal{B}^{\prime}}\right|\geq(1-\epsilon)\left|{\mathcal{B}}\right|$, and $\mathcal{D}$ is a uniformly random $k$-out bipartite multigraph on $\mathcal{A}^{\prime}\times\mathcal{B}^{\prime}$. The algorithm for constructing $\mathcal{D}$ is polynomial time.
Proof.
Fix $\delta=\min\left\{\epsilon/2,(2e^{2}\binom{k+3}{4})^{-1/2}\right\}$. Let $\lambda$ be the value such that the probability that a Poisson random variable $Z\sim\operatorname{Po}(\lambda)$ is less than $k+3$ is $\delta/4$; let us write this as $\Pr(\operatorname{Po}(\lambda)<k+3)=\delta/4$. Choose $C=2.1\lambda$.
Assume that $\operatorname{Po}(2\lambda N)\leq CN$; the probability that this fails to hold is exponentially small in $N$. Randomly orient the first $\operatorname{Po}(2\lambda N)$ edges of $\Psi$, so there are $\operatorname{Po}(\lambda N)$ edges out of each part.
Suppose part $\mathcal{A}$ has $m$ nodes. Choose $t$ to be the power of 2 for which $\frac{1}{2}N/t<m\leq N/t$. The number of outedges $Z(v)$ from a node $v\in\mathcal{A}$ is Poisson with intensity $\lambda N/m\geq\lambda N/(N/t)=t\lambda$. We may thus write $Z(v)$ as $Z(v)=Z_{1}(v)+\cdots+Z_{t}(v)+Z^{\prime}(v)$ where every $Z_{i}(v)\sim\operatorname{Po}(\lambda)$, $Z^{\prime}(v)$ is also Poisson, and all are independent. The event $Z(v)<k+3$ occurs only if every $Z_{i}(v)<k+3$.
Let $X$ be the number of nodes of $\mathcal{A}$ with degree less than $k+3$. From the above, $X=\sum_{v\in\mathcal{A}}\mathbf{1}(Z(v)<k+3)\leq\sum_{v\in\mathcal{A}}{\mathbf{1}}\!\left[{(Z_{1}(v)<k+3)\wedge\cdots\wedge(Z_{t}(v)<k+3)}\right]$, so the event $X\geq\delta m$ requires that, of all $tm$ independent probability-$\delta/4$ events, at least $t\delta m$ hold. Then, $\Pr(X\geq\delta m)\leq\Pr(\operatorname{Bin}(tm,\delta/4)\geq t\delta m)\leq\Pr(\operatorname{Bin}(N,\delta/4)\geq\delta N/2)$. With $\delta$ constant, this has probability exponentially small in $N$.
Then, with exponentially small failure probability, all but a $1-\delta$ fraction of vertices in both $\mathcal{A}$ and (by the same argument, and a union bound) in $\mathcal{B}$ have at least $k+3$ outedges. For any node with more than $k+3$ outedges, randomly delete some to leave exactly $k+3$.
Let $\mathcal{D}_{0}$ be the resulting directed bipartite graph. Let $X_{0}$ be the vertices failing to have outdegree $k+3$. Let $\mathcal{D}$ be the $k$-core of $\mathcal{D}_{0}\setminus X_{0}$. $\mathcal{D}$ can be constructed from $\mathcal{D}_{0}$ by initialising $X=X_{0}$, successively adding to $X$ any node with less than $k$ outneighbours in $\mathcal{D}_{0}\setminus X$, or equivalently with at least $4$ outneighbours in $X$. When there are no more such vertices, $\mathcal{D}$ is the subgraph of $\mathcal{D}_{0}$ induced by $\mathcal{A}^{\prime}=\mathcal{A}\setminus X$ and $\mathcal{B}^{\prime}=\mathcal{B}\setminus X$.
We claim that, at the end, $\left|{X\cap\mathcal{A}}\right|<2\delta\left|{\mathcal{A}}\right|$ and $\left|{X\cap\mathcal{B}}\right|<2\delta\left|{\mathcal{B}}\right|$ a.a.s. By choice of $\delta$, this condition immediately implies the lemma’s conclusion. Suppose the condition does not hold. At any time, let $S$ and $T$ be the vertices added in parts $\mathcal{A}$ and, respectively, $\mathcal{B}$ that bring $X_{0}$ to $X$. (I.e., $S=(X\cap\mathcal{A})\setminus(X_{0}\cap\mathcal{A})$.) For the condition to be violated, at the end we must have $\left|{S}\right|\geq\delta\left|{\mathcal{A}}\right|$ or $\left|{T}\right|\geq\delta\left|{\mathcal{B}}\right|$. Consider the first moment that either event first occurs. There are two symmetric cases, and we consider just the one where, at that moment, $\left|{S}\right|=\delta\left|{\mathcal{A}}\right|$ (up to integrality) while $\left|{T}\right|<\delta\left|{\mathcal{B}}\right|$. Note that every node in $S$ has at least 4 outneighbours in $T\cup(X_{0}\cap\mathcal{B})$. This remains true if we expand $T$ arbitrarily (but disjoint from $X_{0}$) so that $\left|{T}\right|=\delta\left|{\mathcal{B}}\right|$. The probability that any such pair $S,T$ exists is at most the expected number of such pairs, over all fixed subsets and subject to the randomness of $\mathcal{D}$. For a fixed $v\in S$, the probability $v$ has at least $4$ outneighbours in $T$ is at most $\binom{k+3}{4}\left({\left|{T}\right|}/{\left|{\mathcal{B}}\right|}\right)^{4}=\binom{k+3}{4}\delta^{4}$. Then, the expected number of pairs $S,T$ is at most
$\binom{N}{\delta N}\binom{N}{\delta N}\left(\binom{k+3}{4}\delta^{4}\right)^{\delta N}$ $\leq\left(\frac{eN}{\delta N}\right)^{2\cdot\delta N}\left(\binom{k+3}{4}\delta^{4}\right)^{\delta N}$
$=\left(e^{2}\binom{k+3}{4}\delta^{2}\right)^{\delta N}\leq(1/2)^{\delta N}=o(1),$
since we chose $\delta$ to ensure that $e^{2}\binom{k+3}{4}\delta^{2}\leq 1/2$.
So far, we have a graph $\mathcal{D}$ on parts $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ of the claimed cardinalities. Our implementation of the core process revealed only edges directed into $X$, so the edges in $\mathcal{D}$ remain completely random. Each node in $\mathcal{D}$ has outdegree between $k$ and $k+3$. Reveal the outdegree of each node; if a node has outdegree greater than $k$, randomly delete outedges so that it has degree exactly $k$. The remaining edges are still uniformly random, and $\mathcal{D}$ is precisely $k$-out. ∎
Claim 7.
In a uniformly random $k$-out digraph with parts $\mathcal{A}$ and $\mathcal{B}$ both of size $\Theta(N)$, the number of nodes with any double inedge is at most $N^{2/3}$, with failure probability of order $\exp(-\Omega(N^{1/3}))$.
Proof.
We consider double inedges to part $\mathcal{B}$; the symmetric argument applies to part $\mathcal{A}$. Call a node in $\mathcal{B}$ “bad” if it has a multiple inedge. Consider the outedges of a node $a\in\mathcal{A}$ in sequence. If an edge duplicates a previous one (and is the first to do so), it results in a bad $\mathcal{B}$ node. The expected number of such duplications on $a\in\mathcal{A}$ is $O(1/N)$, for a total expected $O(1)$ duplications over all $a\in\mathcal{A}$. Changing an outedge of any $a\in\mathcal{A}$ changes the number of bad $\mathcal{B}$ nodes by at most 1. Azuma’s inequality yields the claim. ∎
Claim 8.
For any constants $k$ and $d$, in a uniformly random $k$-out digraph with parts $\mathcal{A}$ and $\mathcal{B}$ both of size $\Theta(N)$, the probability that $d$ random nodes in $\mathcal{B}$ have a common inneighbour in $\mathcal{A}$ is $O(1/N)$, and the symmetric condition also holds.
Proof.
If $d$ nodes in $\mathcal{B}$ have $a\in\mathcal{A}$ as a common inneighbour, then at least two of them must be in $N^{+}(a)$. Since $\left|{N^{+}(a)}\right|=k$ while $\left|{\mathcal{B}}\right|=\Theta(N)$, this has probability $O(1/N^{2})$. Taking the union bound over all $a\in\mathcal{A}$, the probability the $d$ nodes have any common inneighbour is $O(1/N)$. ∎
Lemma 9 (expansion).
For any $c,k,\delta>0$, let $\eta_{c,k}(\delta)=\max\left\{\frac{17}{3}\delta\ln(e/\delta),2\frac{1-c}{c}k\delta\right\}$. Let $\mathcal{D}$ be a uniformly random $k$-out bipartite multigraph on $\mathcal{A}\times\mathcal{B}$, with $\mathcal{A}\cup\mathcal{B}=[N]$ and $\left|{\mathcal{A}}\right|,\left|{\mathcal{B}}\right|\geq cN$. Then, a.a.s., for every $S\subseteq[N]$ with $\left|{S}\right|\leq\delta N$,
$\left|{N^{-}_{\mathcal{D}}(S)}\right|\leq\sum_{v\in S}\left|{N^{-}_{\mathcal{D}}(v)}\right|\leq\eta_{c,k}(\delta)\,N.$
Note:
  • The lemma refers to all subsets $S$ of size at most $\delta N$ simultaneously, not merely any given subset.
  • We can make $\eta_{c,k}(\delta)$ arbitrarily small by choosing $\delta$ sufficiently small.
  • The lemma implies that sets of size $\delta N$ expand by a factor at most $\eta_{c,k}(\delta)/\delta$. This is false for smaller sets: there will typically be a handful of vertices of arbitrarily high indegree.
  • The proof below uses the first-moment method. A slightly smaller $\eta_{c,k}(\delta)$ can be obtained by choosing (in the Poisson limit) a degree threshold above which there are at least $\delta N$ vertices (thus encompassing the worst-case set $S$), and considering the total of those vertices’ degrees (concentrated, and at least $\left|{N^{-}(S)}\right|$).
Proof.
Assume, without loss of generality, that $\left|{\mathcal{A}}\right|\leq\left|{\mathcal{B}}\right|$. Let $\deg^{-}(v)$ denote the indegree of $v$. For any set $S\in[N]$, define the volume $\operatorname{vol}(S):=\sum_{v\in S}\deg^{-}(v)$ to be the number of edges directed into $S$.
For any set $S$, the neighbourhood size $\left|{N^{-}_{\mathcal{D}}(S)}\right|$ is at most $\operatorname{vol}(S)$; this establishes the lemma’s first inequality. For the second, it is enough for prove the statement for sets $S$ of size $\delta N$.
Let $\rho=\frac{1-c}{c}$. The outneighbours of each node are chosen uniformly with replacement from the other part. So, the indegree of a node $v\in\mathcal{A}$ is distributed as $\operatorname{Bin}(k\left|{\mathcal{B}}\right|,1/\left|{\mathcal{A}}\right|)$, with expectation $k\left|{\mathcal{B}}\right|/\left|{\mathcal{A}}\right|\leq\rho k$. The same bound holds for $v\in\mathcal{B}$ since $\left|{\mathcal{A}}\right|\leq\left|{\mathcal{B}}\right|$. For a given set $S$ of size $\delta N$, $Z=\operatorname{vol}(S)$ is a sum of the two types of binomials, both independent, so $Z$ is dominated by a binomial random variable $Z^{\prime}$ with $\mathbb{E}Z^{\prime}=\rho k\delta N$.
Let $\rho=\frac{1-c}{c}$ and let $\eta=\eta_{c,k}(\delta)$. Apply Chernoff’s inequality $\Pr(Z\geq\mathbb{E}Z+t)\leq\exp\left(-\frac{t^{2}}{2(\mathbb{E}Z+t/3)}\right)$ [15, eq. (2.5)] to $\Pr(Z^{\prime}\geq\eta N)$. This means taking $t=\eta N-\mathbb{E}Z^{\prime}=(r-1)\rho k\delta N$, where $r=\eta/(\rho k\delta)\geq 2$ by definition of $\eta_{c,k}(\delta)$. Then,
$\Pr(Z^{\prime}\geq\eta N)$ $\leq\exp\left(-\frac{(r-1)^{2}(\mathbb{E}Z^{\prime})^{2}}{2(1+\frac{r-1}{3})\mathbb{E}Z^{\prime}}\right)\;\leq\;\exp\left(-\frac{(r/2)^{2}}{2(\frac{2r}{3})}\mathbb{E}Z^{\prime}\right)\;=\;\exp\Big{(}-\tfrac{3}{16}r\cdot\rho\delta kN\Big{)}.$
The number of choices for the set $S$ is
$\binom{N}{\delta N}\leq\left(\frac{eN}{\delta N}\right)^{\delta N}=\exp(\delta N\ln(e/\delta)).$
By the first-moment method, the probability that any set $S$ violates the claimed condition is at most
$\binom{N}{\delta N}\Pr(Z^{\prime}\geq\eta N)$ $\leq\exp\Big{(}\delta N\cdot\Big{(}\ln(e/\delta)-\frac{3}{16}r\rho k\Big{)}\Big{)}$
$=\exp\Big{(}\delta N\cdot\Big{(}\ln(e/\delta)-\frac{3}{16}\frac{\eta}{\delta}\Big{)}\Big{)}=\exp(-\Omega(N))\;=\;o(1),$
the penultimate equality by definition of $\eta_{c,k}(\delta)$. ∎
Lemma 10 (Hall).
Given $k\geq 10$ and $\rho\leq 1/100$. Let $\mathcal{D}$ be a uniformly random $k$-out bipartite digraph on parts $\mathcal{A}$ and $\mathcal{B}$. Consider the subgraph $\mathcal{D}^{\prime}$ induced by a pair of subsets $\mathcal{A}^{\prime}\subseteq\mathcal{A}$, $\mathcal{B}^{\prime}\subseteq\mathcal{B}$, where
  1. (i)
    $\left|{\mathcal{A}^{\prime}}\right|+\left|{\mathcal{B}^{\prime}}\right|\geq(1-\rho)(\left|{\mathcal{A}}\right|+\left|{\mathcal{B}}\right|)$,
  2. (ii)
    $\left|{\mathcal{A}^{\prime}}\right|=\left|{\mathcal{B}^{\prime}}\right|$; and
  3. (iii)
    every node in $\mathcal{D}^{\prime}$ has outdegree at least $k-1$.
Then, a.a.s. in $\left|{\mathcal{A}}\right|$, every such induced subgraph $\mathcal{D}^{\prime}$ contains a perfect matching.
Note that the asymptotic statement in terms of $\left|{\mathcal{A}}\right|$ is equivalent to one in terms of $\left|{\mathcal{B}}\right|$, $\left|{\mathcal{A}^{\prime}}\right|$, or $\left|{\mathcal{B}^{\prime}}\right|$. The hypothesis $k\geq 10$ can be weakened to $k\geq 7$, at the expense of a slightly more complicated proof, and likely further.
Proof.
Let $n:=\left|{\mathcal{A}^{\prime}}\right|=\left|{\mathcal{B}^{\prime}}\right|$. If $\mathcal{D}^{\prime}$ does not have a perfect matching (ignoring the edge directions) then there are Hall sets $S\subseteq\mathcal{A}^{\prime}$ and $T\subseteq\mathcal{B}^{\prime}$ such that $N(S)\subseteq T$ and $\left|{T}\right|=\left|{S}\right|-1$. Write $\bar{S}$ for $\mathcal{A}^{\prime}\setminus S$ and $\bar{T}$ for $\mathcal{B}^{\prime}\setminus T$. If $S,T$ is a Hall pair, so is $\bar{T},\bar{S}$. Since $\left|{S}\right|+\left|{\bar{T}}\right|=n+1$, at least one of them must be at most $(n+1)/2$. By symmetry, it suffices to consider a Hall pair $S,T$ where $S\subseteq\mathcal{A}$ has cardinality $s\leq(n+1)/2$.
Henceforth we ignore $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$. We are simply interested in the existence of a Hall pair $S,T$ with $S\subseteq\mathcal{A}$, $T\subseteq\mathcal{B}$, $\left|{S}\right|=s\leq(n+1)/2$, and $\left|{T}\right|=s-1$.The probability that $\mathcal{D}$ allows such a pair is at most the expected number of such pairs, which is at most the number of pairs times the probability that a fixed pair has the Hall property.
Since $2n=\left|{\mathcal{A}^{\prime}}\right|+\left|{\mathcal{B}^{\prime}}\right|\geq(1-\rho)(\left|{\mathcal{A}}\right|+\left|{\mathcal{B}}\right|)$, we have that $\left|{\mathcal{A}}\right|+\left|{\mathcal{B}}\right|\leq\frac{2}{1-\rho}n$ and thus
$\left|{\mathcal{A}}\right|$ $\leq\tfrac{2}{1-\rho}n-\left|{\mathcal{B}}\right|\;\leq\;\tfrac{2}{1-\rho}n-\left|{\mathcal{B}^{\prime}}\right|\;=\;\left(\tfrac{2}{1-\rho}-1\right)n\;\leq\;\tfrac{3}{e}n;$
the last inequality follows from the hypothesis on $\rho$ and is used in the next calculation. The same holds for $\mathcal{B}^{\prime}$. Also, $\left|{\mathcal{A}}\right|$ and $\left|{\mathcal{B}}\right|$ are both $\Theta(n)$, and henceforth we work asymptotically in $n$.
For each $v\in S$, all but one of $v$’s outedges must be directed into $T$. The probability that there is such a Hall pair $S,T$, then, is at most the number of choices of $S\subseteq\mathcal{A}$ with $\left|{S}\right|=s$, of $T\subseteq\mathcal{B}$ with $\left|{T}\right|=s-1$, and (by hypothesis iii) of one edge per element of $S$, times the probability that such a combination has all other edges of $S$ directed into $T$. The probability there is any such pair is at most
$\binom{k}{1}^{s}\binom{(3/e)n}{s}\binom{(3/e)n}{s-1}\left(\frac{s-1}{n}\right)^{(k-1)s}$ $\leq k^{s}\left(\frac{3n}{s}\right)^{s}\left(\frac{3n}{s-1}\right)^{s-1}\left(\frac{s-1}{n}\right)^{(k-1)s}$
$\leq(9k)^{s}\left(\frac{s-1}{n}\right)^{(k-3)s+1}$
$\leq\frac{s}{n}\left(9k\left(\frac{s-1}{n}\right)^{k-3}\right)^{s}.$ (3.1)
[…]
4. Matching Strategy and Analysis
We construct a hypergraph $H$, initially empty and eventually containing a perfect matching $\mathcal{M}$, on vertex set $V=[n]$; we assume that $n$ is divisible by 3. Partition $V$ into a set $A_{0}$ of vertices called “apexes”, and a set $B_{0}$ of pairs of vertices, called “base pairs” or just “bases”, with
$\left|{A_{0}}\right|$ $=\frac{n}{3}+4\epsilon_{D}n$ $\left|{B_{0}}\right|$ $=\frac{n}{3}-2\epsilon_{D}n,$ (4.1)
where $\epsilon_{D}$ is a constant chosen small enough to satisfy various claims below. The two members of a base pair are sometimes called “partners”. Integrality is not an issue; all that needs to be exact is that $\left|{A_{0}}\right|+2\left|{B_{0}}\right|=n$.
4.1. Phase 1: A robust matching structure
In this phase we construct the digraph $D_{1}$ introduced above, and in tandem the corresponding hypergraph $H$.
For a sequence of semirandom offers, do the following. For each offer of the form $\left\{a,b\right\}$, where $a\in A_{0}$ and $b\in\mathbf{b}\in B_{0}$ is a member of a base pair, choose as third point the partner of $b$. This defines a hyperedge $\left\{a\right\}\cup\mathbf{b}$, which we add to our hypergraph $H$, and an (undirected) auxiliary graph edge $\left\{a,\mathbf{b}\right\}$, which we add to a set $\Psi$. Semi-random offers not of the specified form are ignored. (Recall what this means from the beginning of Section 2.)
Note that each $(a,\mathbf{b})\in\Psi$ is uniformly random in $A_{0}\times B_{0}$, with replacement. Given $\epsilon_{D}$ and $k$, let $C$ be the corresponding value in Lemma 6. Each semirandom offer is of the desired form w.p. (with probability) $\frac{1}{3}\cdot\frac{2}{3}+\frac{2}{3}\cdot\frac{1}{3}+O(\epsilon_{D})=4/9+O(\epsilon_{D})>1/3$ for $\epsilon_{D}$ sufficiently small, so in $3Cn$ semirandom offers, a.a.s. there are at least $Cn$ of the stipulated form, exceeding the lemma’s $CN$ (since by (4.1), $N$ is about $\frac{2}{3}n$).
Should this a.a.s. statement fail, we consider the entire construction to have failed. We will take the same approach throughout, proceeding on the assumption that all the a.a.s. statements hold.
It follows from Lemma 6 that from this set $\Psi$ of undirected edges we can a.a.s. construct a bipartite digraph $D_{1}$ on parts $A_{1}\subseteq A_{0}$ and $B_{1}\subseteq B_{0}$, with $\left|{A_{1}}\right|\geq(1-\epsilon_{D})\left|{A_{0}}\right|$ and $\left|{B_{1}}\right|\geq(1-\epsilon_{D})\left|{B_{0}}\right|$, and $D_{1}$ is uniformly $k$-out on these node sets.
By construction, $H$ contains the hyperedges corresponding to the edges of $D_{1}$. We will discard (or ignore) all other hyperedges in $H$ so that there is a one-to-one correspondence between the edges in $D_{1}$ and the hyperedges in $H$: they are two representations of the same thing.
Let $X_{1}=(A_{0}\setminus A_{1})\cup(B_{0}\setminus B_{1})$ be the set of “failed” nodes missing from $D_{1}$. By Lemma 6 and (4.1), we have
$\left|{A_{1}}\right|$ $\geq(1-\epsilon_{D})\left|{A_{0}}\right|>\frac{n}{3}+3\epsilon_{D}n$ (4.2)
$\left|{B_{1}}\right|$ $\geq(1-\epsilon_{D})\left|{B_{0}}\right|>\frac{n}{3}-3\epsilon_{D}n$ (4.3)
$\left|{X_{1}}\right|$ $\leq\epsilon_{D}(\left|{A_{0}}\right|+\left|{B_{0}}\right|)<\epsilon_{D}n.$ (4.4)
Let $\dot{X}_{1}$ be the set of all the points (hypergraph vertices) contained in $X_{1}$, so $\dot{X}_{1}=(A_{0}\cap X_{1})\cup\bigcup_{\mathbf{b}\in(B_{0}\cap X_{1})}\mathbf{b}$. Letting $a_{x}=\left|{A_{0}\cap X_{1}}\right|$ and $b_{x}=\left|{B_{0}\cap X_{1}}\right|$, note that
$|{\dot{X}_{1}}|$ $=a_{x}+2b_{x}.$ (4.5)
4.2. Phase 2: Matching exceptional points
This phase will place every point of $\dot{X}_{1}$ into a hyperedge to be included in the matching $\mathcal{M}$, while keeping $D$ robust enough that it will contain a matching.
At the beginning of this phase, $A=A_{1}$, $B=B_{1}$, $D=D_{1}$, $X=X_{1}$, and $\dot{X}=\dot{X}_{1}$. Along the way, $A,B,D,X,\dot{X}$ will change; e.g., points will be removed from $\dot{X}$ as they are placed in hyperedges of $\mathcal{M}$.
We first consider the operations we will perform in this phase, deferring the question of how many semirandom offers are required.
4.2.1. Phase 2a
Here we treat all the points in $X_{1}$, from apexes and base pairs alike. We will be more precise in a moment, but roughly speaking, any time we get a semirandom offer consisting of two points $a,a^{\prime}\in A$, we choose an arbitrary $x\in\dot{X}$, defining a hyperedge $e=\left\{a,a^{\prime},x\right\}$, and add $e$ irrevocably to $\mathcal{M}$. The failed vertex $x$ is now in a hyperedge in $\mathcal{M}$, but $a$ and $a^{\prime}$ must be deleted from $A$ and $D$: neither is any longer available to be combined with a base pair.
As said in the Outline (Section 3), we will ensure that, as we delete nodes from $D$, every node in $D$ retains outdegree at least $k-1$. To do this, we “block” some nodes of $D$ from future deletion; we let $Q$ denote the set of blocked nodes. To avoid that any single node’s deletion decreases an inneighbour’s outdegree by 2 or more, we intialise $Q$ to be the set of nodes with parallel inedges. To avoid that the outdegree of any node $u$ is decreased more than once, if ever an outneighbour $v$ of $u$ is deleted, we block all other outneighbours of $u$ from future deletion. That is, whenever a node $v$ is deleted from $D$, we add $N^{+}(N^{-}(v))$ to $Q$. All that remains is to ensure that in any single action, we do not delete two nodes with a common inneighbour.
With this out of the way, we can be precise. For any semirandom offer consisting of two points $a,a^{\prime}\in A_{0}$, we choose an arbitrary $x\in\dot{X}$, defining a hyperedge $e=\left\{a,a^{\prime},x\right\}$. We discard $e$ unless the following conditions both hold:
  1. (M2a-C1)
    $a,a^{\prime}\in A\setminus Q$.
  2. (M2a-C2)
    $N^{-}(a)$ and $N^{-}(a^{\prime})$ are disjoint.
(Read the label as M for matching, 2a for the phase, and C for “condition”, and in the next part, A for “action”.)
If the conditions hold, we take the following action:
  1. (M2a-A1)
    Add $e$, irrevocably, to $\mathcal{M}$.
  2. (M2a-A2)
    Delete $x$ from $\dot{X}$ and $a$ and $a^{\prime}$ from $A$ (and hence also from $D$).
  3. (M2a-A3)
    Add $N^{+}(N^{-}(\left\{a,a^{\prime}\right\}))$ to $Q$, blocking the outneighbours of the inneighbours of the deleted nodes.
Each such action resolves one point in $\dot{X}$, so by (4.5), $a_{x}+2b_{x}$ actions are required. Each action deletes 2 nodes of part $A$ and 0 nodes of part $B$, so at the end of this phase, $\left|{B}\right|=b_{0}-b_{x}$ and $\left|{A}\right|=(a_{0}-a_{x})-2(a_{x}+2b_{x})=a_{0}-3a_{x}-4b_{x}$. Also,
$\left|{A}\right|-\left|{B}\right|$ $=a_{0}-b_{0}-3a_{x}-3b_{x}\;=\;a_{0}-b_{0}-3\left|{X_{1}}\right|$ (4.6)
[…]
Thus, $A$ is larger than $B$ at the end of Phase 2a.
4.2.2. Phase 2b
This phase makes the sizes of $A$ and $B$ exactly equal so that they can be matched in $D$. Recall that at the end of Phase 2a, $\left|{A}\right|>\left|{B}\right|$. To even them up, we will generate hyperedges each consisting of three vertices from $A$, deleting the points from $A$ and adding the hyperedge to $\mathcal{M}$.
To be precise, each time we get a semirandom offer consisting of two vertices $a,a^{\prime}\in A_{0}$, we define a hyperedge $e=\left\{a,a^{\prime},a^{\prime\prime}\right\}$ for a uniformly random $a^{\prime\prime}\in A\setminus\left\{a,a^{\prime}\right\}$. We discard the offer unless the following conditions both hold:
  1. (M2b-C1)
    $a,a^{\prime},a^{\prime\prime}\in A\setminus Q$.
  2. (M2b-C2)
    $N^{-}(a)$, $N^{-}(a^{\prime})$, and $N^{-}(a^{\prime\prime})$ are disjoint.
If the conditions hold, we take the following action:
  1. (M2b-A1)
    Add $e$, irrevocably, to $\mathcal{M}$.
  2. (M2b-A2)
    Delete $a,a^{\prime}$, and $a^{\prime\prime}$ from $A$ (and hence also from $D$).
  3. (M2b-A3)
    Add $N^{+}(N^{-}(\left\{a,a^{\prime},a^{\prime\prime}\right\}))$ to $Q$, blocking the outneighbours of the inneighbours of the deleted nodes.
From (4.6), at the beginning of Phase 2b we have $\left|{A}\right|-\left|{B}\right|\equiv a_{0}-b_{0}\equiv a_{0}+2b_{0}=n\equiv 0\pmod{3}$. Since each action deletes 3 vertices from $A$, we can make $\left|{A}\right|=\left|{B}\right|$ exactly. The precise number of actions needed is not critical — it is obviously $O(\epsilon_{D}n)$ — but from (4.6) and (4.1) it is
$\frac{\left|{A}\right|-\left|{B}\right|}{3}=\frac{a_{0}-b_{0}}{3}-\left|{X_{1}}\right|\leq\frac{a_{0}-b_{0}}{3}=2\epsilon_{D}n.$ (4.8)
4.2.3. Analysis
How many nodes can ever be blocked? Initially, $Q=Q_{1}$ consists of the nodes with double inedges. By 7, $|{Q_{1}}|=O(n^{2/3})$, with failure probability exponentially small in $n$. As usual, we will declare failure of the construction if this fails.
Recall that when an action deletes a node $v$ from $D$, it adds $N^{+}(N^{-}(v))$ to $Q$. Phase 2a takes $a_{x}+2b_{x}$ actions, each deleting 2 nodes from $D$; note that $a_{x}+2b_{x}\leq 2|{X_{1}}|\leq 2\epsilon_{D}n$, by (4.4). Phase 2b takes at most $2\epsilon_{D}n$ actions, each deleting 3 nodes from $D$. Altogether, they delete at most $10\epsilon_{D}n$ nodes from $D$.
The number of inneighbours of these nodes is bounded by Lemma 9. The lemma’s $N$ is $\left|{A_{1}}\right|+\left|{B_{1}}\right|>\frac{2}{3}n$ by (4.2) and (4.3), so $10\epsilon_{D}n$ represents at most a $15\epsilon_{D}$ fraction of all nodes; we take the lemma’s $\delta$ equal to $15\epsilon_{D}$. The two parts $A_{1}$ and $B_{1}$ are of nearly equal size, so we take the lemma’s $c$ as $\frac{1}{3}$. By the lemma, the number of inneighbours is at most $\eta_{1/3,k}(\epsilon_{D})N<\eta_{1/3,k}(\epsilon_{D})n$.
In the paragraph above, we reasoned that the parts are each of size about $\frac{1}{3}n$, and that (with generous allowance for the small difference) each is of size at least $c=1/3$ their combined size. There will be several such cases, so let us treat this one more carefully as an example. Each size is $\frac{1}{3}n+O(\epsilon_{D})n$, where the $O(\cdot)$ term indicates a value, positive or negative, bounded in absolute value by some constant such as 15 that we could work out by tracing calculations such as (4.1) and (4.2). We omit these calculations and similar ones, as we will take $\epsilon_{D}$ small enough that these $O(\cdot)$ terms are tiny, much less than the generous margins we will allow. More to the point, forgetting the constants calculated here, it is clear that $\epsilon_{D}$ can be chosen sufficiently small to make it so.
Each inneighbour has at most $k$ outneighbours, so the number of outneighbours is at most $k\>\eta_{1/3,k}(\epsilon_{D})n$. We would like this always to be small compared with $\left|{A}\right|$ and $\left|{B}\right|$, both of which will always be very nearly $\frac{1}{3}n$, and safely at least $\frac{1}{4}n$.
This dictates our parameter choices. We will contrive that $k\>\eta_{1/3,k}(\epsilon_{D})\leq\frac{1}{101}\cdot\frac{1}{4}$, ensuring that the size of $Q$, including $Q_{1}$, satisfies $|{Q}|\leq\frac{1}{100}\min\left\{\left|{A_{0}}\right|,\left|{B_{0}}\right|\right\}$ at all times. We take $k=10$ for purposes of Lemma 10. We find that $\delta=$0.000\hskip 1.30005pt015$$ suffices, giving $\eta_{10,1/3}(\delta)<$0.001$<1/404$ as desired. (The value of $\eta$, defined in Lemma 9, is determined by the first term in the max; the second term, $40\>\delta$, is much smaller.) Since we set $\delta=15\epsilon_{D}$, taking $\epsilon_{D}=$0.000\hskip 1.30005pt001$$ suffices.
Looking ahead, we will invoke Lemma 10 to prove existence of a perfect matching in $D$. That hypothesises that, in the current context, the final set $\mathcal{A}$ has size $\left|{\mathcal{A}}\right|\geq(1-\rho)\left|{\mathcal{A}_{1}}\right|$, and likewise for $\mathcal{B}$, with $\rho=1/100$. Since the difference between $\mathcal{A}_{1}$ and $\mathcal{A}$ is just the $O(\epsilon_{D})n$ nodes deleted in Phase 2, this hypothesis is clearly satisfied, again without bothering to work out the small constant hidden in the $O(\cdot)$.
Having controlled the size of $Q$ we are ready to consider the number of semirandom offers needed in Phase 2. In Phase 2a, that the first point offered is in $A_{0}$ occurs with probability just above $1/3$. The cardinality of $A$ is nearly that of $A_{0}$ and we have arranged that $\left|{Q}\right|\leq\frac{1}{100}\left|{A_{0}}\right|$, so $a\in A\setminus Q$ with probability at least
$4\displaystyle 1/3\cdot(1-O(\epsilon_{Q})-1/100)>1/$ 4 . (4.9)
Independently (almost), the same holds for $a^{\prime}$, so (M2a-C1) holds with probability at least 1/161161/161 / 16.111We would have independence if the offered $a^{\prime}$ were independent of $a$, allowing the possibility that $a^{\prime}=a$ (in which case we would reject the offer). We can simulate this independence by tossing a coin and with probability $1/n$ replacing the offered $a^{\prime}$ with $a$. In the simulated model, (M2a-C1) and distinctness hold with probability at least $1/16-1/n$, or simply $1/16$ by absorbing the $1/n$ into slack in (4.9). Since the simulated substitution of $a^{\prime}$ for $a$ is never advantageous, (M2a-C1) holds with probability at least $1/16$ in the original model.
By 8, the chance that (M2a-C2) fails to hold is $O(1/n)$. Thus, each offer satisfies the conditions with probability at least $1/17$. Since at most $2\epsilon_{D}n$ actions are needed (see (4.4)), a.a.s., $18\cdot 2\epsilon_{D}n$ offers suffice.
Similarly, in Phase 2b, an offer plus the added point $a^{\prime\prime}$ satisfies condition (M2b-C1) with probability more than $(1/4)^{3}$ (even if we chose $a^{\prime\prime}$ in [$n$] rather than $A$), and again (M2b-C2) fails with probability $O(1/n)$, so each offer results in action with probability at least $1/65$. By (4.8), the number of actions needed is at most $2\epsilon_{D}n$, so, a.a.s., $131\epsilon_{D}n$ offers suffice.
Numerically, Phases 2a and 2b together require far under $0.001$n$ offers. More to the point, it is clear that, with control of $\left|{Q}\right|$, each offer results in action with probability $\Theta(1)$, and thus $O(n)$ offers suffice.
Recall that Phase 1 used $3Cn$ semirandom offers, with $C$ given by Lemma 6. Now that we have fixed $\epsilon_{D}=$0.000\hskip 1.30005pt001$$, the proof of Lemma 6 requires a $\lambda$ of 41, for $C$ a bit under $100$, so $300n$ semirandom offers suffice for Phase 1 (and the whole algorithm). We have made no effort whatsoever to optimise the constants.
We use $A_{2},B_{2},D_{2}$ to denote $A,B,D$ at the end of Phase 2. At this point, all points outside of $A_{2},B_{2}$ appear in hyperedges of $\mathcal{M}$. Since no nodes were deleted from $B$ during Phase 2, and applying (4.3), we have:
$\left|{A_{2}}\right|$ $=\left|{B_{2}}\right|=\left|{B_{1}}\right|>\frac{n}{3}-3\epsilon_{D}n.$ (4.10)
4.3. Phase 3: Matching the bulk of the points
We complete the hypergraph matching $\mathcal{M}$ with hyperedges corresponding to the edges of a perfect matching in $D_{2}$; Lemma 10 proves that a.a.s. $D_{2}$ has a perfect matching.
We apply the lemma with $\mathcal{D}=D_{1}$ and $\mathcal{D}^{\prime}=D_{2}$, and accordingly $\mathcal{A}=A_{1}$, $\mathcal{A}^{\prime}=A_{2}$, and $\mathcal{B}=\mathcal{B}^{\prime}=B_{1}=B_{2}$.
We already noted in Section 4.2.3 (and confirmed in (4.10)) that hypothesis i is satisfied: the fractional size difference between $A_{1}$ and $A_{2}$ is $O(\epsilon_{D})$ and far less than $1/100$. Hypothesis ii, that the two parts are of equal cardinality, is satisfied by construction; see just before (4.8). Hypothesis iii, that every node degree is $k$ or $k-1$, is also satisfied by construction, specifically the various action conditions based on the blocked set $Q$. Thus, a.a.s. $D_{2}$ contains a perfect matching.
This concludes the proof of Theorem 1.
[…]
We now establish Theorem 2. As already explained, constructing a perfect matching in the 1-offer $s$-uniform model can trivially be simulated in the 2-offer $s$-uniform model, simulating a 2nd “offer” vertex with the extra “chosen” vertex.
Constructing a matching in the 2-offer $s$-uniform model cannot be done using the 2-offer 3-uniform model as a black box, but can be done in exactly the same way. In short, in Phase 0, take a bit more than $n/s$ vertices as apexes $A_{0}$, arbitrarily partitioning the rest into a bit fewer than $n/s$ “base” groups each of size $s-1$, comprising $B_{0}$.
4.4.1. Phase 1
In Phase 1, when offered an apex $a$ and a point $b$ from a base $\mathbf{b}$, propose the hyperedge $a\cup\mathbf{b}$ and correspondingly an auxiliary-graph edge $\left\{a,\mathbf{b}\right\}$, directed one way or the other. As before, with such proposals we can construct a directed auxiliary graph $D=D_{1}$, bipartite with parts $A_{1}$ and $B_{1}$ comprising most of $A_{0}$ and $B_{0}$ respectively, with $D_{1}$ a uniformly random $k$-out digraph on $A_{1}$ and $B_{1}$. Let $X=X_{1}=(A_{0}\cup B_{0})\setminus(A_{1}\cup B_{1})$ be the nodes missing from $D$, and $\dot{X}$ the set of points in $X$.
4.4.2. Phase 2
In Phase 2a, until $\dot{X}$ is empty, when offered two points in $A_{0}$, propose a hyperedge consisting of them and any $s-2$ points in $\dot{X}$. If in a final step $20<|{\dot{X}}|<s-$ 2, supplement the points of $\dot{X}$ with distinct unblocked points from $A$. Correspondingly, propose the action of deleting the two apexes from part $A$ of the auxiliary graph $D$.
As we did for matchings, we block $N^{+}(N^{-}(v))$ for any node $v$ deleted from $D$. We accept a proposed action (and accept the corresponding hyperedge for $\mathcal{M}$) if the offered vertices lie in the current set $A$ and are not blocked. By analogy with the argument around (4.9), by controlling the size of $\dot{X}_{1}$ we can ensure that the number of nodes deleted and blocked is small, ensuring in turn that each proposed action is accepted w.p. at least $(1/s\cdot 0.98)^{s-2}=\Omega(1)$. Also as for matchings, we choose the sizes $\left|{A_{0}}\right|$ and $\left|{B_{0}}\right|$ so that, a.a.s., when Phase 2a ends, $\left|{A}\right|\geq\left|{B}\right|$.
In Phase 2b, when offered 2 vertices $a,a^{\prime}\in A_{0}$, propose a hyperedge consisting of them and $s-2$ distinct points in $A\setminus\left\{a,a^{\prime}\right\}$ and correspondingly propose the action of deleting these $s$ points from part $A$ of $D$. Accept the proposal if the two offered points were in $A\setminus Q$; in this case, delete these $s$ apexes from $D$, and accept the hyperedge for $\mathcal{M}$.
Again, each proposed action is accepted w.p. at least $(1/s\cdot 0.98)^{s-2}=\Omega(1)$. Stop when the number of apexes is equal to the number of bases, $\left|{A}\right|=\left|{B}\right|$.
For the stopping condition, since at the end of Phase 2a, $\left|{A}\right|\geq\left|{B}\right|$, and Phase 2b decreases the size of $A$ without changing $B$, it is clear that at some time we get $\left|{A}\right|\leq\left|{B}\right|$. At this time, we get equality if $\left|{A}\right|\equiv\left|{B}\right|\pmod{s}$, and this is so. Let $V=[n]$ be the set of all points, with $\left|{V}\right|\equiv 0\pmod{s}$. Let $\dot{V(D)}$ be the set of points in $D$. At this step, the points of $\dot{V(D)}$ are all in hyperedges in $\mathcal{M}$, so $\left|{V\setminus\dot{V(D)}}\right|\equiv 0\pmod{s}$. Thus, at this step,
$\left|{A}\right|-\left|{B}\right|$ $\equiv\left|{A}\right|+(s-1)\left|{B}\right|=\left|{\dot{V(D)}}\right|=\left|{V}\right|-\left|{V\setminus\dot{V(D)}}\right|\equiv 0\pmod{s}.$ (4.11)
4.4.3. Phase 3
Finally, construct a matching in $D$ (ignoring the edge directions). The corresponding hypergraph edges complete a perfect matching in the hypergraph.
As for matchings, it is enough to ensure that when constructing $D_{1}$, $\left|{X}\right|\leq\epsilon_{D}n$ for some sufficiently small constant $\epsilon_{D}$. The number of Phase 2 steps is $O(\epsilon_{D}n)$ (the $O(\cdot)$ expression hiding some constant independent of $\epsilon_{D}$), so the number of nodes deleted from $D$ is also $O(\epsilon_{D}n)$. By Lemma 9, $Q$ remains of size at most $k\>\eta_{c,k}(\delta)(O(\epsilon_{D}))n$ (plus $O(n^{2/3})$), and this can be made an arbitrarily small fraction of $n$ by choosing $\epsilon_{D}$ sufficiently small. If follows that each action is accepted w.p. $\Omega(1)$.
5. Loose Hamilton cycles
Now we turn to the proof of Theorem 3. We will demonstrate a strategy that, within a linear number of steps, a.a.s. produces a 3-uniform hypergraph with a loose Hamilton cycle. The vertex set is $[n]$ for some even $n$. (The number of vertices must be even if a loose Hamilton cycle exists.) The odd points will occur twice each in the loose cycle, the even points just once, in the form
$(1,a_{1},3),\;(3,a_{2},5),\;\ldots,\;(n-1,a_{n/2},1),$ (5.1)
where we write each edge as an ordered triple to indicate the roles of the three points, and $a_{1},a_{2},\ldots,a_{n/2}$ is a permutation of $2,4,\ldots,n$. We will not be able to get exactly such a cycle, with the odd numbers appearing in order, but will obtain a Hamilton cycle that differs from it on very few edges.
The approach is similar to that for generating a perfect matching in Section 4. We define apex and base sets, each of cardinality exactly $n/2$:
$A_{0}$ =$=\left\{2,4,\ldots,n\right\}$ $B_{0}$ =$=\left\{\left\{1,3\right\},\left\{3,5\right\},\ldots\left\{n-1,1\right\}\right\}.$ (5.2)
Here the bases are overlapping, in order to build a cycle like that of (5.1), where for hypergraph matchings they were disjoint.
The next steps are structurally the same as for matchings in Section 4. We will construct a directed bipartite graph on $A_{0}$ and $B_{0}$, but it will not have a perfect matching because a small number of nodes fail to have sufficiently high degree (and cause other nodes to “fail” as well). We deal with the failed nodes first, and in doing so add certain hyperedges, irrevocably, to our cycle-in-the-making $\mathcal{C}$. ($\mathcal{C}$ plays the role that $\mathcal{M}$ did for hypergraph matchings.) When the failed nodes are dealt with, most of our bipartite graph will remain, and what remains will have a perfect matching. The matching edges correspond to hyperedges in $\mathcal{H}$, and they complete $\mathcal{C}$ to a loose Hamilton cycle.
5.1. Phase 1
In the first phase, we build a graph $D$ in the same manner as we did when building a perfect matching in Section 4.
For each offer of the form $\left\{a,b\right\}$, where $a\in A$ and $b\in\mathbf{b}\in B$ is a member of a base pair, choose as third point the partner of $b$ and add $\left\{a,\mathbf{b}\right\}$ to a set $\Psi$. (We ignore any offers that are not of this form.)
$\Psi$ forms a set of edges of a bipartite graph on $A_{0}\times B_{0}$. Note that a perfect matching on this graph would yield a loose Hamilton cycle of the form in (5.1).
For any $\epsilon_{D}>0$, Lemma 6 ensures, exactly as in Phase 1 of Section 4, that within $O(n)$ steps we can construct a directed bipartite graph $D=D_{1}$ with parts $A_{1}\subseteq A_{0}$ and $B_{1}\subseteq B_{0}$, such that $D_{1}$ is a uniformly random $k$-out bipartite multigraph on $A_{1}\times B_{1}$, and a.a.s. $\left|{A_{1}}\right|\geq(1-\epsilon_{D})\left|{A_{0}}\right|$ and $\left|{B_{1}}\right|\geq(1-\epsilon_{D})\left|{B_{0}}\right|$. We refer to the apexes in $A_{1}$ as good and those in $A_{X}=A_{0}\setminus A_{1}$ as failed, and likewise for bases. We have
$a_{x}$ $=\left|{A_{X}}\right|\leq\epsilon_{D}n/2$ $b_{x}$ $=\left|{B_{X}}\right|\leq\epsilon_{D}n/2.$ (5.3)
5.2. Phase 2
We introduce an additional structure, $\mathcal{P}$, central to the analysis. $\mathcal{P}$ is a graph on the odd points, consisting of vertex-disjoint paths. A path in $\mathcal{P}$ can be an isolated vertex. Initially, the edges in $\mathcal{P}$ are precisely the base pairs $B_{1}$. Eventually we will make $\mathcal{P}$ a single path and then a Hamilton cycle. Always, each edge in $\mathcal{P}$ is one of two types: it consists of the two odd points in a hyperedge $e=(b_{1},a,b_{2})\in\mathcal{H}$ (and all the apexes used for such edges will be distinct); or, as in the initial $\mathcal{P}$, it consists of a base pair in $B$ (and later it will be possible to perfectly match all these base pairs and the remaining apexes, in $D$). When $\mathcal{P}$ is a Hamilton cycle, this ensures that there is a corresponding loose hypergraph cycle $\mathcal{C}$ as desired.
In this phase we will again use a set $Q$ of nodes of $D$ “blocked” from use. Initially we will set $Q$ to consist of any nodes whose inedges include any double edge. Initially, $\mathcal{P}$ has $b_{x}$ components. Let us describe the phase just enough to count the number of nodes it will delete from each part of $D$, both to bound $\left|{Q}\right|$ and to confirm that, going in to Phase 3, $\left|{A}\right|=\left|{B}\right|$.
In Phase 2a, each action (see Figure 1) will assign an apex in $A_{X}$ to a hyperedge (along with two base points from different base pairs) and place it in $\mathcal{C}$, increasing the number of paths in $\mathcal{P}$ by 1, and deleting 2 base pairs from $B$. Phase 2a takes $a_{x}$ actions to resolve all the failed apexes, resulting in a total of $2a_{x}$ deletions from the set of base pairs in $B$, so that $\left|{B_{0}\setminus B}\right|=b_{x}+2a_{x}$, and $b_{x}+a_{x}$ components in $\mathcal{P}$. The part $A$ is unchanged, with $\left|{A_{0}\setminus A}\right|=a_{x}$.
In Phase 2b, each action will decrease the number of components of $\mathcal{P}$ by 1 (see Figure 2), deleting 2 apexes from $D$ (committing them to hyperedges) and deleting 1 base pair from $B$. This phase takes $a_{x}+b_{x}$ actions to make $\mathcal{P}$ a cycle, in the process making $\left|{A_{0}\setminus A}\right|=a_{x}+2(a_{x}+b_{x})$ and $\left|{B_{0}\setminus B}\right|=(b_{x}+2a_{x})+(a_{x}+b_{x})$, both equal to $3a_{x}+2b_{x}$.
In the end, then, we have equal numbers of apexes and base pairs not in $D$, and thus also equal numbers remaining in $D$.
The number of apex nodes deleted from $D$ by the two phases is $0+2(a_{x}+b_{x})=2a_{x}+2b_{x}$, and the number of base nodes deleted is $2a_{x}+(a_{x}+b_{x})=3a_{x}+b_{x}$; by (5.3) each is at most $2\epsilon_{D}n$.
Let $S$ be the set of deleted nodes. For any given $\epsilon_{Q}>0$, let Lemma 9’s $\eta$ equal $\epsilon_{Q}/k$, and take the lemma’s corresponding $\delta$ to determine $2\epsilon_{D}$. Then, Lemma 9 ensures that the deleted set’s inneighbourhood has size $\left|{N^{-}(S)}\right|\leq(\epsilon_{Q}/k)n$. Since $D$ is $k$-out, the inneighbourhood’s outneighbourhood — which is to say, $Q$ — has size
$\left|{Q}\right|=\left|{N^{+}(N^{-}(S))}\right|\leq k(\epsilon_{Q}/k)n=\epsilon_{Q}n.$
5.2.1. Phase 2a
First we treat the failed apex nodes, those in $A_{X}$, doing something analogous to Phase 2a for matchings. In Phase 2b we will treat the failed bases, which has some extra complexity.
Upon offer of odd points $b_{1}$ and $b_{2}$, we propose a hyperedge $e=(b_{1},a,b_{2})$, where $a$ is an arbitrary apex in $A_{X}$. If $b_{1}$ and $b_{2}$ belong to a common path $P\in\mathcal{P}$ then, on the induced path between them, let $b^{\prime}_{1}$ be the neighbour of $b_{1}$ and $b^{\prime}_{2}$ that of $b_{2}$. If not, let $b^{\prime}_{1}=b_{1}+2$ and $b^{\prime}_{2}=b_{2}+2$.
We discard the hyperedge unless the following conditions all hold:
  1. (H2a-C1)
    $(b_{1},b_{2})$ is not in $\mathcal{P}$ (therefore not in $B$).
  2. (H2a-C2)
    $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ are in $B\setminus Q$.
  3. (H2a-C3)
    $N^{-}(b_{1})\cap N^{-}(b_{2})=\emptyset$.
(The labelling here is H to connote Hamilton cycle, 2a the phase, and C a condition, with A for action in the next group.) Several of these conditions are unnecessary, but it is probable that they all hold (as will be shown),
$s_{1}$$b_{1}$$b_{1}^{\prime}$$t_{1}$$s_{2}$$b_{2}$$b_{2}^{\prime}$$t_{2}$$(b_{1},a,b_{2})$
Figure 1. Salvaging a “failed” apex $a$. Upon offer of suitable odd points $b_{1}$ and $b_{2}$, hyperedge $(b_{1},a,b_{2})$ is introduced (dashed line) and added to $\mathcal{C}$. The good base pairs $\left\{b_{1},b_{1}^{\prime}\right\}$ and $\left\{b_{2},b_{2}^{\prime}\right\}$ (dotted lines) are deleted from $D$. (As a failed apex, $a$ was already absent from $D$.) This converts two paths in $\mathcal{P}$ (the figure’s upper and lower lines) into three, or one path into two if $b_{1}$ and $b_{2}$ lie on a common path (if in the figure, $t_{1}=t_{2}$).
If the conditions all hold, take the following action (see Figure 1).
  1. (H2a-A1)
    Add $(b_{1},a,b_{2})$ irrevocably to $\mathcal{C}$, and add $(b_{1},b_{2})$ to $\mathcal{P}$.
  2. (H2a-A2)
    Delete $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ from $\mathcal{P}$, $B$, and $D$. Also, delete $a$ from $A_{X}$.
  3. (H2a-A3)
    Add $N^{+}(N^{-}(\left\{(b_{1},b^{\prime}_{1}),(b_{2},b^{\prime}_{2})\right\}))$ to $Q$, blocking the outneighbours of the inneighbours of the deleted nodes.
The action incorporates the failed apex $a$ into a hyperedge, at the expense of removing two good base pairs $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ from $D$, and turning two paths in $\mathcal{P}$ into three (as shown in the figure, when $b_{1}$ and $b_{2}$ are in distinct paths) or one into two (when $b_{1}$ and $b_{2}$ are in a common path). The action decreases the outdegree of any node in $D$ by at most 1: it is only $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ that are deleted, by condition (H2a-C2) neither was in $Q$ therefore neither has a double inedge, and by condition (H2a-C3) they have no common inneighbour.
In any trial, the conditions are likely to be satisfied. For (H2a-C1), there is probability $O(1/n)$ that $b_{1}$ and $b_{2}$ are path neighbours. For (H2a-C2), $\mathcal{P}$ has $(1/2-O(\epsilon_{D}))n$ edges on its $n/2$ points, so most points have two edges in $\mathcal{P}$ (so it is likely $b_{1}$ and $b_{2}$ do) while only $O(\epsilon_{Q})n$ points are in $Q$ (so it is unlikely that $b_{1}$ or $b_{2}$ are). Finally, (H2a-C3) follows from 8. Since all of the handful of failure events have probability $O(\epsilon_{D}+\epsilon_{Q})$, so does their union.
5.2.2. Phase 2b
We now turn to the $O(\epsilon_{D})n$ failed bases, which has some extra complexity. Details will follow but the basic idea is illustrated in Figure 2. If there were a previous semirandom offer of $(b,a)$ accepted as hyperedge $(b,a,s_{1})$, with $s_{1}$ an endpoint of a path $P_{1}\in\mathcal{P}$ whose other endpoint is $t_{1}$, then if there is a new offer $(b^{\prime},a^{\prime})$, where $(b,b^{\prime})$ is an edge in a path $P$ in $\mathcal{P}$, accepting the new offer as $(b^{\prime},a^{\prime},t_{1})$ allows the paths $P$ and $P_{1}$ to be merged. In this case, the hyperedges $(b,a,s_{1})$ and $(b^{\prime},a^{\prime},t_{1})$ are both added irrevocably to $\mathcal{C}$ and $(b,b^{\prime})$ is deleted from $P$, $B$, and $D$.
The operation works equally well whether $b^{\prime}$ lies between $b$ and $t$, as shown, or between $b$ and $s$. If $\mathcal{P}$ consists of a single path, $s,\ldots,b,b^{\prime},\ldots,t$, then a similar operation, using hyperedges $(b,a,t)$ and $(b^{\prime},a^{\prime},s)$, turns $\mathcal{P}$ into a cycle.
$s$$b$$b^{\prime}$$t$$s_{1}$$t_{1}$$(b,a,s_{1})$$(b^{\prime},a^{\prime},t_{1})$
Figure 2. Patching together base-pair paths. Suppose a previous offer included an apex $a$ and base point $b$, to which we added some path end $s_{1}$ to propose hyperedge $(b,a,s_{1})$. If by coincidence the current offer includes base point $b+2$ and an apex $a^{\prime}$, then, for some the opposite path end $t_{1}$, propose hyperedge $(b+2,a^{\prime},t_{1})$. Accepting this pair of proposals means adding the (dashed) path edges $\left\{b,s_{1}\right\}$ and $\left\{b+2,t_{1}\right\}$ and deleting the (dotted) edge $\left\{b,b+2\right\}$, turning two $\mathcal{P}$ paths into one. Correspondingly, we accept the two hyperedges into $\mathcal{C}$, and delete the base pair $\left\{b,b+2\right\}$ from $B$.
This merging phase proceeds in rounds, each starting with $\mathcal{P}$ having $\ell$ components and reducing that to $\left\lfloor(9/10)\ell\right\rfloor$ components, until the final round where $\mathcal{P}$ goes from a Hamilton path to a Hamilton cycle. If any round fails, we declare failure of the whole algorithm; we will show this to be unlikely.
We partition the $n/2$ odd points into $n/4$ “cell” pairs, $\left\{1,3\right\},\allowbreak\left\{5,7\right\},\allowbreak\ldots,\allowbreak\left\{n-3,n-1\right\}$; these pairs are fixed for all rounds. Hyperedges added in a round are used only within that round; except for those used in actions, they are ignored in future rounds and can be thought of as deleted.
Algorithm for a round
We consider a round starting when $\mathcal{P}$ has $\ell$ components. For each path $P_{i}$ in $\mathcal{P}$, designate one endpoint as the start $s_{i}$ and the other as the terminal $t_{i}$. (If $P_{i}$ is an isolated point, $s_{i}=t_{i}$.) The round will consider up to $10f$ semirandom offers, with
$f$ $=f(\ell)=n^{3/4}\ell^{1/4}.$ (5.4)
The round terminates with success as soon as $\mathcal{P}$ has $\left\lfloor(9/10)\ell\right\rfloor$ components, or with failure after $10f$ semirandom offers or after $f^{2}/n$ actions have been attempted (see below), whichever comes first.
Consider only offers of the form $(b,a)$, odd and even, discarding others. Also discard any offer where $b$ was offered earlier in this round. Assume then that this is the first time $b$ is offered. If the cell partner $b^{\prime}$ of $b$ also has not been offered, complete this offer to $(b,a,t)$ where $t$ is a random terminal, and add $(b,a,t)$ to $\mathcal{H}$. Consider the cell “seeded”. If the cell partner $b^{\prime}$ was previously seen (the cell was seeded), then some hyperedge $(b^{\prime},a^{\prime},t_{i})$ is in $\mathcal{H}$, in which case complete the current offer to $(b,a,s_{i})$. Consider the cell of $b$ “filled”, and “attempt action”. Attempting action means taking action (see below) if the following conditions are satisfied:
  1. (H2b-C1)
    The path $P_{i}$ with endpoints $s_{i}$ and $t_{i}$ has not already been merged into another path: $s_{i}$ and $t_{i}$ remain path endpoints. ($P_{i}$ may have subsumed another path; that is fine.)
  2. (H2b-C2)
    $(b,b^{\prime})\notin P_{i}$. (This condition does not apply in the final round; see below.)
  3. (H2b-C3)
    $(b,b^{\prime})$ is a good base pair in $B$ (and therefore an edge in $\mathcal{P}$).
  4. (H2b-C4)
    $a$ and $a^{\prime}$ are distinct apexes in $A$.
  5. (H2b-C5)
    None of $(b,b^{\prime})$, $a$, nor $a^{\prime}$ is in $Q$ (blocked).
  6. (H2b-C6)
    $N^{-}(a)\cap N^{-}(a^{\prime})=\emptyset$.
If these conditions are satisfied, the action is taken as follows. The action decreases the number of components of $\mathcal{P}$ by 1; see Figure 2.
  1. (H2b-A1)
    Add the hyperedges $(b,a,s_{i})$ and $(b^{\prime},a^{\prime},t_{i})$ to $\mathcal{C}$, and add the edges $(b,s_{i})$ and $(b^{\prime},t_{i})$ to $\mathcal{P}$.
  2. (H2b-A2)
    Delete $(b,b^{\prime})$ from $\mathcal{P}$, $B$, and $D$, and delete $a$ and $a^{\prime}$ from $A$ and $D$.
  3. (H2b-A3)
    Add $N^{+}(N^{-}(\left\{(b,b^{\prime}),a,a^{\prime}\right\}))$ to $Q$, blocking the outneighbours of the inneighbours of the deleted nodes.
The last round is a special case. Any round starting with $\ell\geq 10$ paths ends with $\left\lfloor\frac{9}{10}\ell\right\rfloor\geq 9$. The rounds starting with $2\leq\ell\leq 9$ paths each decrease the number of paths by 1, so the last round starts with $\ell=1$. Here, if $b$ and $b^{\prime}$ share a cell in the single path $s,\ldots,b,b^{\prime},\ldots,t$, then if offered $(b,a)$ we make hyperedge $(b,a,t)$, and if offered $(b^{\prime},a^{\prime})$ we make hyperedge $(b^{\prime},a^{\prime},s)$. In this case, an action makes $\mathcal{P}$ a Hamilton cycle. At that point, we proceed to Phase 3.
Analysis
The conditions ensure that no node in $D$ has its outdegree reduced by more than 1.
Also, Phase 2b completes within $O(n)$ offers. Round 0 starts with $\ell_{0}=O(n)$ so round $i$ starts with $\ell_{i}\leq\ell_{0}(9/10)^{i}$, and over all rounds the total number of offers is at most
$\sum 10f(\ell_{i})$ $=\sum 10n^{3/4}((9/10)^{i}\ell_{0})^{1/4}=O(n)\sum((9/10)^{1/4})^{i}=O(n).$
It remains only to show that a.a.s., in every round, at least $(1/10)\ell$ actions are taken.
Note that for every $i$, with $f_{i}=f(\ell_{i})$,
$f_{i}$ $\geq f(1)=\Omega(n^{3/4})$ and $f_{i}$ $\leq f(\ell_{0})=n\epsilon_{D}^{1/4}\leq n/100,$ (5.5)
the latter depending on having $\epsilon_{D}$ sufficiently small.
We now focus on a single round and thus refer simply to $\ell$ and $f=f(\ell)$. We will use the phrase “overwhelmingly unlikely” for probabilities of order $\exp(-\Omega(n^{p}))$ for some constant $p>0$, and “with overwhelming probability” for the complement.
Claim 11.
With probability $1-\exp(-\Omega(n^{1/4}))$, $10f$ offers fill at least $f^{2}/n$ cells, enabling $f^{2}/n$ attempted actions.
Proof.
Although the offers arrive successively (and the round is terminated once $(1/10)\ell$ actions are taken), consider the round’s full $10f$ offers. Each offer has probability just over $1/2$ of being of the requisite form $(b,a)$. Each one of that form has equal probability of $b$ being a “left” or “right” element of a cell, i.e., of $b\equiv 1\pmod{4}$ or $b\equiv 3\pmod{4}$. In $10f$ offers, with probability $1-\exp(-\Omega(n))$, there will be at least $2.4f$ each lefts and rights; we assume henceforth that this is so.
We have $\ell\leq\ell_{0}\leq O(\epsilon_{D})n$ and thus (for $\epsilon_{D}$ sufficiently small) $102.4f(\ell)\leq n/$ 10 and thus each new left falls into a new cell with probability at least $9/10$, so the expected number of left-occupied cells is at least $2.2f$. Each of these left “balls” fell in a random cell, and changing the location of one ball changes the number of left-occupied cells by at most 1, so by the Azuma-Hoeffding inequality the probability of a deviation of the order of the mean is overwhelmingly unlikely: of order
$\exp(-\Omega(f^{2}/(f\cdot 1)))=\exp(-\Omega(f)).$
We presume henceforth that there are at least $2.1f$ left-occupied cells.
Likewise, the right balls are overwhelmingly likely to fall into $2.1f$ distinct cells, and we presume henceforth that they do. The locations of these right-occupied cells are uniformly random. Each has probability at least $(2.1f)/(n/4)$ of also being left-occupied, so the expected number of filled cells is at least $17.6f^{2}/n$. Changing the location of one right-occupied cell changes the number of full cells by at most 1, so by the Azuma-Hoeffding inequality the probability of a deviation on the order of the mean is overwhelmingly unlikely, of order
$\exp(-\Omega((f^{2}/n)^{2}/(f\cdot 1)))$ $=\exp(-\Omega(f^{3}/n^{2}))=\exp(-\Omega(n^{1/4})).$
Claim 12.
When a new cell is filled, the conditions for action are satisfied with probability at least $2/10$, assuming that the number of paths in $\mathcal{P}$ is at most $(9/10)\ell$.
Proof.
The location of the new cell is uniformly random over all cells, conditioned on not being one of the cells previously filled. For a uniformly random cell $X$, let $F$ be the event that $X$ was not previously filled, and let $E$ be the event that $X$ satisfies the conditions for action. We are interested in $\Pr(E\mid F)$; consider first just $\Pr(E)$.
For a random cell $X$, condition (H2b-C1) fails w.p. at most $1/10$, or the round would already have ended successfully. Condition (H2b-C2) fails only if the randomly selected terminal is the terminal of the path containing $b$, and the probability of this is $21/\ell\leq 1/$ 2. Conditions (H2b-C3) and (H2b-C4) fail with probability $O(\epsilon_{D})$, because $A$ and $B$ contain all but an $O(\epsilon_{D})$ fraction of $A_{0}$ and $B_{0}$ respectively. Condition (H2b-C5) fails w.p. $O(\epsilon_{Q})$ because $\left|{Q}\right|=O(\epsilon_{Q})n$ throughout. Finally, (H2b-C6) fails w.p. $O(1/n)$ by 8. In all,
$\Pr(E)$ $\geq 1-(1/10+1/2+O(\epsilon_{D}+\epsilon_{Q})+O(1/n))\geq 3/10$
for appropriate choices of $\epsilon_{D}$ and $\epsilon_{Q}$.
The algorithm terminates if the number of attempted actions — cells filled — exceeds $f^{2}/n$. Thus, the probability that a random cell $X$ was previously filled is $\Pr(\overline{F})\leq(f^{2}/n)/(n/4)=4(f/n)^{2}\leq 2/100^{2}$ by (5.5). So, $\Pr(F)\geq 0.99$. It follows that $\Pr(E\mid F)\geq\Pr(E)/\Pr(F)\geq 2/10$. ∎
From 11, with overwhelming probability at least $f^{2}/n$ actions can be attempted, and if the hypothesis of 12 is satisfied, the number of these succeeding is distributed as $B(f^{2}/n,2/10)$, and thus is overwhelmingly likely to exceed $\frac{1}{10}f^{2}/n$: using (5.5) again, the failure probability is of order $\exp(-\Omega(f^{2}/n))=\exp(-\Omega(n^{1/2}))$.
This number of successful actions would far exceed $\ell$: their ratio is
$\frac{\frac{1}{2}f^{2}/n}{\ell}$ $=\frac{n^{3/2}\ell^{1/2}}{2n\ell}=\frac{1}{2}(n/\ell)^{1/2}\geq\frac{1}{2}(1/\epsilon_{D})^{1/2}>1.$
It is impossible that the number of successful actions exceeds $\ell$, so we conclude that, with overwhelming probability, the hypothesis of 12 must at some point fail. That is, at some point $\mathcal{P}$ must fall below $(9/10)\ell$, upon which the round terminates with success.
Since the failures were all overwhelmingly unlikely (of order $\exp(-\Omega(n^{p}))$ for some constant $p>0$), by the union bound, failure remains overwhelmingly unlikely even over the $O(\ln n)$ rounds.
Assuming success, Phase 2 terminates with $\mathcal{P}$ consisting of a single cycle. Some of its edges correspond to hyperedges committed to $\mathcal{C}$; its other edges consist of base pairs comprising one part of $D$, whose other part consists of the apexes not yet committed to $\mathcal{C}$. Thus $D$ is an induced subgraph of $D_{1}$, in which every node’s degree is at most 1 smaller than it is in $D_{1}$.
5.3. Phase 3
As previously stated, in this final phase we take a perfect matching between $A$ and $B$ in the graph $D$ with directions removed from its edges. And, as noted earlier, the cardinalities of $A$ and $B$ match. Lemma 10 implies that this matching exists, just as in Subsection 4.3.
For each $\{b,b+2\}\in B$ and its matching partner $a\in A$, we place the hyperedge $\{b,a,b+2\}$ into $\mathcal{C}$. That hyperedge was formed in Phase 1, when the corresponding edge was added to $D$. That completes the Hamilton cycle.
[…]
The linear-time 2-offer strategy to construct a loose Hamilton cycle in a 3-uniform hypergraph extends easily to $s$-uniform hypergraphs.
To avoid excessive notation, consider $s=4$ for example. Take base pairs $\left\{1,4\right\},\left\{4,7\right\},\left\{7,10\right\},\ldots$, and apex nodes (no longer single points) $\left\{2,3\right\},\left\{5,6\right\},\left\{8,9\right\},\ldots$. The generalisation to other $s$ is clear. We will see that it is easy to keep each apex node set intact, so that it behaves just like an apex point in the $s=3$ case.
In Phase 1, on offer of a base point $b$ and an apex point $a$ contained in an apex set $\mathbf{a}\in A$, choose base point $b+s-1$ (to give base pair $\mathbf{b}=\left\{b,b+s-1\right\}$), choose all the other points in $\mathbf{a}$ (to give apex $\mathbf{a}$), creating an edge between $\mathbf{b}$ and $\mathbf{a}$ in $D$ and correspondingly creating the hyperedge $\left\{\mathbf{b}\cup\mathbf{a}\right\}$. This gives an auxiliary graph $D$ as before.
In Phase 2a, on offer of base points $\left\{b_{1},b_{2}\right\}$, make hyperedge $\left\{\left\{b_{1},b_{2}\right\}\cup\mathbf{a}\right\}$ where $\mathbf{a}$ is a failed apex node. As before, this uses up the node $\mathbf{a}$ and turns two paths into three.
In Phase 2b, for an offered base point $b$ and apex point $a\in\mathbf{a}\in A$, choose some path end point $s_{1}$ and the rest of $\mathbf{a}$ to make candidate hyperedge $\left\{\left\{b,s_{1}\right\}\cup\mathbf{a}\right\}$. When another such candidate hyperedge contains $b+s-1$, just as for $s=3$, three paths can be turned into two, using these two hyperedges and deleting the base $\left\{b,b+s-1\right\}$.
With these extensions, the mechanics is exactly as for the $s=3$ case. The treatment of the auxiliary graph $D$ is identical. This gives the strategy for constructing a loose Hamilton cycle in the 2-offer model for an $s$-uniform hypergraph.
Of course, whatever can be done in linearly many steps in the 2-offer model can also be done in the 1-offer model. This establishes Theorem 4.
Acknowledgements
Greg is grateful to Alan Frieze for discussions of Hamilton cycles, and to Olaf Parczyk for extensive discussions of the semirandom hypergraph process.